How many liters of H, gas at STP can be produced


by the reaction of 4.60 g of Na and excess water

1
Expert's answer
2022-04-22T04:50:05-0400

Solution:

The molar mass of Na is 23 g/mol

Therefore,

Moles of Na = (4.60 g Na) × (1 mol Na / 23 g Na) = 0.20 mol Na


Balanced chemical equation:

2Na + 2H2O → 2NaOH + H2

According to stoichiometry:

2 mol of Na produce 1 mol of H2

Thus, 0.20 mol of Na produce:

(0.20 mol Na) × (1 mol H2 / 2 mol Na) = 0.1 mol H2


At STP, one mole of any gas occupies a volume of 22.4 L.

Thus, 0.1 mol of H2 occupies:

(0.1 mol H2) × (22.4 L / 1 mol) = 2.24 L H2

Volume of H2 = 2.24 L


Answer: 2.24 liters of H2